# A note on the oriented line graph operation

### v2

@@ Abstract
Oriented line graph, introduced by Kotani and Sunada (2000), is closely related to Hashimato’s non-backtracking matrix (1989). It is known that for regular graphs $G$, the eigenvalues of the adjacency matrix of the oriented line graph $\vec{L}(G)$ of $G$ are the reciprocals of the poles of the Ihara zeta function of $G$. We determine the characteristic polynomials of the adjacency matrix of the underlying undirected graph of $\vec{L}(G)$ and the skew-symmetric adjacency matrix (and Hermitian adjacency matrix) of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$.
A locally bijective (resp. injective) homomorphism from a graph $G$ to a graph $H$ is a mapping $\psi\colon V(G)\to V(H)$ such that for every vertex $v$ of $G$, the restriction of $\psi$ to the neighborhood $N_{G}(v)$ is a bijection (resp. injection) from $N_{G}(v)$ to $N_{H}(\psi(v))$ (Fiala and Kratochvíl, 2008). An out-neighborhood bijective (resp. injective) homomorphism from a directed graph $\vec{G}$ to a directed graph $\vec{H}$ is a mapping $\psi\colon V(\vec{G})\to V(\vec{H})$ such that for every vertex $v$ of $\vec{G}$, the restriction of $\psi$ to the out-neighborhood $N_{\vec{G}}^{+}(v)$ is a bijection (resp. injection) from $N_{\vec{G}}^{+}(v)$ to $N_{\vec{H}}^{+}(\psi(v))$ (Antony and Shalu, 2025). We prove that the existence of a locally bijective (resp. injective) homomorphism from a graph $G$ of minimum degree at least 3 to a graph $H$ is equivalent to the existence of an out-neighborhood bijective (resp. injective) homomorphism from $\vec{L}(G)$ to $\vec{L}(H)$. We also prove some results on the coloring variants distance-two coloring and star coloring.
@@ 1 Introduction
The notion of oriented line graphs is introduced by Kotani and Sunada [11], and they are closely related to Hashimato’s [10] non-backtracking matrix. Let $G$ be a (simple undirected) graph. The oriented line graph of $G$ is the oriented graph with vertex set $\bigcup_{uv\in E(G)}\{(u,v),(v,u)\}$ and there is an arc from a vertex $(u,v)$ to a vertex $(v,w)$ in it when $u\neq w$ [17]. We denote the oriented line graph of $G$ by $\vec{L}(G)$, and its underlying undirected graph by $L^{*}(G)$. We can rephrase the definition of oriented line graph in terms of $\vec{D}(G)$, the simple directed graph obtained from $G$ be replacing each edge $uv$ of $G$ by two arcs $(u,v)$ and $(v,u)$ (i.e., replacing each edge by a digon). The oriented line graph of $G$ is the oriented graph with the set of arcs of $\vec{D}(G)$ as its vertex set, and there is an arc in it from a vertex $(u,v)$ to a vertex $(v,w)$ when $u\neq w$ (see Fig. 1). The non-backtracking matrix of $G$ is the adjacency matrix of $\vec{L}(G)$.
@@
The characteristic polynomial of the adjacency matrix of $\vec{L}(G)$ is well-known for $d$-regular graphs $G$ with $d\geq 3$. It is known that the roots of this polynomial are the reciprocals of the poles of the Ihara zeta function of $G$ [10]. We determine the characteristic polynomial of (the adjacency matrix of) $L^{*}(G)$ and the skew-symmetric adjacency matrix of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$. We prove that the existence of certain types of locally constrained homomorphisms from a graph $G$ to a graph $H$ is equivalent to the existence of certain types of locally constrained homomorphisms from $\vec{L}(G)$ to $\vec{L}(H)$. We also prove some results on the coloring variants distance-two coloring and star coloring.
@@
This paper is organized as follows. Section 2 discusses our results related to the spectrum of some matrices associated with the oriented line graph operation. Section 3 discusses the results involving graph homomorphisms, orientations and colorings. Section 3 includes the subsections 3.1 giving necessary definitions, 3.2 dealing with graph homomorphisms, and 3.3 dealing with distance-two coloring and star coloring. We conclude with Section 4.
2 Spectrum of $L^{*}(G)$
In this section, we determine the characteristic polynomial of (the adjacency matrix of) $L^{*}(G)$, and the skew-symmetric adjacency matrix and Hermitian adjacency matrix of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$.
Lubetzky and Peres [15] proved that the adjacency matrix of an oriented line graph is unitarily similar to a block-diagonal matrix with $1\times 1$ and $2\times 2$ blocks.
Theorem 2.1(Lubetzky and Peres [15]).
@@ 2 Spectrum of $L^{*}(G)$
where 1-1 and $1$ are repeated mnm-n times,
@@
Using Theorem 2.1, we determine the characteristic polynomial of $L^{*}(G)$.
@@
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the characteristic polynomial of $L^{*}(G)$ is $(x^{2}-4)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})^{2}-(d-2)^{2}\right)$.
@@
Note that $\theta_{i}+\theta_{i}^{\prime}=\lambda_{i}$, and $\theta_{i}\theta_{i}^{\prime}=d-1$ for all $i$. Let $B$ denote non-backtracking matrix of $G$; that is, the adjacency matrix of $\vec{L}(G)$. By Theorem 2.1, $B$ is unitarily similar to the matrix $C$ of Equation 1. That is, $B=P^{*}CP$ for some unitary matrix $P$ (i.e., $P^{-1}=P^{*}$). Note that $B+B^{t}$ is the adjacency matrix of $L^{*}(G)$. Also, $B+B^{t}=B+B^{*}=P^{*}CP+(P^{*}CP)^{*}=P^{*}CP+P^{*}C^{*}P=P^{*}(C+C^{*})P$. That is, $B+B^{t}$ is unitarily similar to $C+C^{*}$.
@@
where 1-1 and $1$ (resp. 2-2 and $2$) are repeated mnm-n times. If $|\lambda_{i}|\leq 2\sqrt{d-1}$, then $\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu=\lambda_{i}=\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu$. If $|\lambda_{i}|>2\sqrt{d-1}$, then $\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu=2\theta_{i}$, and $\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu=2\theta_{i}^{\prime}$.

### v3

# Some short notes on oriented line graphs and related matrices

@@ Abstract
Oriented line graph, introduced by Kotani and Sunada (2000), is closely related to Hashimato’s non-backtracking matrix (1989). It is known that for regular graphs $G$, the eigenvalues of the adjacency matrix of the oriented line graph $\vec{L}(G)$ of $G$ are the reciprocals of the poles of the Ihara zeta function of $G$. We determine the characteristic polynomial of the $z$-Hermitian adjacency matrix of $\vec{L}(G)$ for each $z\in\mathbb{C}$ and $d$-regular graph $G$ with $d\geq 3$. Special cases of this matrix include the Hermitian adjacency matrix of $\vec{L}(G)$ and the adjacency matrix of the underlying undirected graph of $\vec{L}(G)$. We also exhibit an application to star coloring of graphs.
@@ 1 Introduction
The notion of oriented line graphs is introduced by Kotani and Sunada [kotani_sunada], and they are closely related to Hashimato’s [hashimoto] non-backtracking matrix. Let $G$ be a (simple undirected) graph. The oriented line graph of $G$ is the oriented graph with vertex set $\bigcup_{uv\in E(G)}\{(u,v),(v,u)\}$ and there is an arc from a vertex $(u,v)$ to a vertex $(v,w)$ in it when $u\neq w$ [parzanchevski]. We denote the oriented line graph of $G$ by $\vec{L}(G)$, and its underlying undirected graph by $L^{*}(G)$. We can rephrase the definition of oriented line graph in terms of $\vec{D}(G)$, the simple directed graph obtained from $G$ be replacing each edge uvuv of $G$ by two arcs $(u,v)$ and $(v,u)$ (i.e., replacing each edge by a digon). The oriented line graph of $G$ is the oriented graph with the set of arcs of $\vec{D}(G)$ as its vertex set, and there is an arc in it from a vertex $(u,v)$ to a vertex $(v,w)$ when $u\neq w$ (see Fig. 1). The non-backtracking matrix of $G$ is the adjacency matrix of $\vec{L}(G)$.
@@
For $z\in\mathbb{C}$ and an oriented graph $\vec{H}$ with vertex set $\{v_{1},v_{2},\dots,v_{n}\}$, the $z$-Hermitian adjacency matrix of $\vec{H}$ is the matrix whose $(i,j)$-th entry is 11 if $(v_{i},v_{j}),(v_{j},v_{i})\in E(\vec{H})$; $z$ if $(v_{i},v_{j})\in E(\vec{H})$ but $(v_{j},v_{i})\notin E(\vec{H})$; $\bar{z}$ if $(v_{j},v_{i})\in E(\vec{H})$ but $(v_{i},v_{j})\notin E(\vec{H})$; and 0 if $(v_{i},v_{j}),(v_{j},v_{i})\notin E(\vec{H})$.
The characteristic polynomial of the adjacency matrix of $\vec{L}(G)$ is well-known for $d$-regular graphs $G$ with $d\geq 3$. It is known that the roots of this polynomial are the reciprocals of the poles of the Ihara zeta function of $G$ [hashimoto]. We determine the characteristic polynomial of the $z$-Hermitian adjacency matrix of $L^{*}(G)$ for each $z\in\mathbb{C}$ and $d$-regular graph $G$ with $d\geq 3$. Special cases of this matrix include the Hermitian adjacency matrix of $\vec{L}(G)$ and the adjacency matrix of $L^{*}(G)$. It is interesting that for a $d$-regular graph $G$ with $d\geq 3$, the line graph has the characteristic polynomial $\operatorname{char}(L(G);x)=(x+2)^{m-n}\prod_{i=1}^{n}(x-\lambda_{i}-d+2)$, whereas $L^{*}(G)$ has the characteristic polynomial $\operatorname{char}(L^{*}(G);x)=(x+2)^{m-n}(x-2)^{m-n}\prod_{i=1}^{n}(x-\lambda_{i}-d+2)(x-\lambda_{i}+d-2)$. We also exhibit an application to star coloring of graphs (see Section 3 for definition).
@@
[…]
We determine the characteristic polynomial of the $z$-Hermitian adjacency matrix of $\vec{L}(G)$ for each $z\in\mathbb{C}$ and $d$-regular graph $G$ with $d\geq 3$.
Lubetzky and Peres [lubetzky_peres] proved that the adjacency matrix of an oriented line graph is unitarily similar to a block-diagonal matrix with $1\times 1$ and $2\times 2$ blocks.
Theorem 2.1(Lubetzky and Peres [lubetzky_peres]).
@@
where 1-1 and 11 are repeated mnm-n times,
@@
Using Theorem 2.1, we determine the $z$-Hermitian spectrum of $\vec{L}(G)$.
Theorem $2.$
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues
$\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the $z$-Hermitian characteristic polynomial of $\vec{L}(G)$ where $z\in\mathbb{C}$ is
$(x^{2}-(z+\bar{z})^{2})^{m-n}\prod_{i=1}^{n}\left[(x-z\lambda_{i})(x-\bar{z}\lambda_{i})-|z|^{2}d^{2}+(z+\bar{z})^{2}(d-1)\right]$
Proof overview.
Let $B$ be the non-backtracking matrix of $G$. Then $B$ is unitarily similar to the block diagonal matrix $C$ in Theorem 2.1. Thus $\theta_{j}+\theta_{j}^{\prime}=\lambda_{j}$,  $\theta_{j}\theta_{j}^{\prime}=(d-1)$, and $\theta_{j}^{2}+\theta_{j}^{\prime 2}=\lambda_{j}^{2}-2(d-1)$.
[…]
$\left.\begin{bmatrix}z\theta_{n}+\bar{z}\bar{\theta}_{n}&z\alpha_{n}\\ \bar{z}\bar{\alpha}_{n}&z\theta_{n}^{\prime}+\bar{z}\bar{\theta}_{n}^{\prime}\end{bmatrix},-z-\bar{z},\dotsc,-z-\bar{z},z+\bar{z},\dotsc,z+\bar{z}\right)$.
For $\lambda_{j}\leq 2\sqrt{d-1}$, $|\alpha_{j}|^{2}=(d-2)^{2}$ and $\theta_{j}^{\prime}=\bar{\theta}_{j}$. Therefore,
$(x^{2}-(z+\bar{z})^{2})^{m-n}\prod_{i=1}^{n}\left((x-z\lambda_{i})(x-\bar{z}\lambda_{i})-|z|^{2}d^{2}+(z+\bar{z})^{2}(d-1)\right)$
For $\lambda_{j}>2\sqrt{d-1}$, $|\alpha_{j}|^{2}=d^{2}-\lambda_{j}^{2}$ and $\bar{\theta}_{j}=\theta_{j}$. Therefore,
$(x^{2}-(z+\bar{z})^{2})^{m-n}\prod_{i=1}^{n}\left((x-z\lambda_{i})(x-\bar{z}\lambda_{i})-|z|^{2}d^{2}+(z+\bar{z})^{2}(d-1)\right)$
Corollary 1.
The $\omega$-Hermitian characteristic polynomial of $\vec{L}(G)$ is
$(x^{2}-1)^{m-n}\prod_{i=1}^{n}\left[(x^{2}+x\lambda_{i}+\lambda_{i}^{2})-(d^{2}-d+1)\right]\qquad$
Corollary 2.
The characteristic polynomial of $L^{\ast}(G)$ is
$(x^{2}-4)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})^{2}-(d-2)^{2}\right)$
A direct proof of Corollary 2 is provided in Appendix.
Corollary 3.
$\operatorname{char}(L(\vec{D}(G));x)=x^{m-n}\prod_{i=1}^{n}(x-\lambda_{i})$ [beineke_bagga2021a] (note that $L(\vec{H})$ denotes the line digraph of a digraph $\vec{H}$).
@@
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues
$\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the characteristic polynomial of the skew-symmetric adjacency matrix of $\vec{L}(G)$ is $x^{2(m-n)}\prod_{i=1}^{n}(x^{2}+d^{2}-\lambda_{i}^{2})$. ∎
The next theorem follows from Lemma 3.1 of Bilu and Linial [bilu_linial] since $L^{*}(G)$ is a 2-lift of $L(G)$ [cyriac_shalu1], and the characteristic polynomials of $L^{*}(G)$ and $L(G)$ are $(x^{2}-4)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})^{2}-(d-2)^{2}\right)$ and $(x+2)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})-(d-2)\right)$,
respectively.
Theorem 2.4.
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues
$\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Let $\psi$ be an LBH from $L^{*}(G)$ to $L(G)$. Let $\{V_{0},V_{1}\}$ be a partition of the vertex set of $\vec{L}(G)$ such that $|\psi^{-1}(v)\cap V_{i}|=1$ for each vertex $v$ of $G$ and each $i\in\mathbb{Z}_{2}$. Let $(L(G),s)$ be the signed graph such that for each edge uvuv of $L(G)$, we have $s(uv)=1$ if and only if uvuv belongs to the cut $(V_{0},V_{1})$ in $\vec{L}(G)$. Then, the characteristic polynomial of the signed adjacency matrix of $(L(G),s)$ is $(x-2)^{m-n}\prod_{i=1}^{n}(x-\lambda_{i}+d-2)$. ∎
3 Consequence to star coloring
For a positive integer $q$, a star $q$-coloring of a graph $G$ is a function $f\colon V(G)\to\mathbb{Z}_{q}$ such that (i) $f(u)\neq f(v)$ for every edeg uvuv of $G$, and (ii) there is no path $u,v,w,x$ in $G$ with $f(u)=f(w)$ and $f(v)=f(x)$. A Locally Bijective Homomorphism (LBH) from a graph $G$ to a graph $H$ is a mapping $\psi\colon V(G)\to V(H)$ such that for every vertex $v$ of $G$, the restriction of $\psi$ to the neighbourhood $N_{G}(v)$ is a bijection from $N_{G}(v)$ onto $N_{H}(\psi(v))$ [fiala_kratochvil].
For $p\geq 2$, a $K_{1,p+1}$-free $2p$-regular graph $G$ is star $(p+2)$-colorable if and only if $G$ admits an LBH to $L^{*}(K_{p+2})$ [cyriac_shalu1]. Since existence of LBH from a graph $H$ to a graph $J$ implies that the characteristic polynomial of $J$ divides the characteristic polynomial of $H$ [fiala_kratochvil], we have the following by Corollary 2.
Theorem 3.1.
Let $G$ be a $K_{1,p+1}$-free $2p$-regular star $(p+2)$-colorable graph, where $p\geq 2$. Then, the characteristic polynomial of ((the adjacency matrix of)) $G$ is divisible by $(x-2p)(x+2)^{(p-1)(p+2)/2}$ $(x-2)^{p(p+1)/2}(x-p+2)^{p+1}(x+p)^{p+1}$. ∎
[…]
Appendix 0.A Direct proof for spectrum of $L^{*}(G)$
Theorem 0.A.1.
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues
$\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the characteristic polynomial of $L^{*}(G)$ is exactly $(x^{2}-4)^{m-n}$
$\prod_{i=1}^{n}\left((x-\lambda_{i})^{2}-(d-2)^{2}\right)$.
@@ Appendix 0.A Direct proof for spectrum of $L^{*}(G)$
Note that $\theta_{i}+\theta_{i}^{\prime}=\lambda_{i}$, and $\theta_{i}\theta_{i}^{\prime}=d-1$ for all ii. Let $B$ denote non-backtracking matrix of $G$; that is, the adjacency matrix of $\vec{L}(G)$. By Theorem 2.1, $B$ is unitarily similar to the matrix $C$ of Equation 1. That is, $B=P^{*}CP$ for some unitary matrix $P$ (i.e., $P^{-1}=P^{*}$). Note that $B+B^{t}$ is the adjacency matrix of $L^{*}(G)$. Also, $B+B^{t}=B+B^{*}=P^{*}CP+(P^{*}CP)^{*}=P^{*}CP+P^{*}C^{*}P=P^{*}(C+C^{*})P$. That is, $B+B^{t}$ is unitarily similar to $C+C^{*}$.
@@
where 1-1 and 11 (resp. 2-2 and 22) are repeated mnm-n times. If $|\lambda_{i}|\leq 2\sqrt{d-1}$, then $\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu=\lambda_{i}=\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu$. If $|\lambda_{i}|>2\sqrt{d-1}$, then $\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu=2\theta_{i}$, and $\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu=2\theta_{i}^{\prime}$.